OpenStudy (anonymous):

Solve the equation: 4x2 = 9

OpenStudy (anonymous):

Answers: {-2, -3} {2/3, -2/3} {3/2, -3/2} no solution

OpenStudy (anonymous):

I would say {3/2, -3/2}

OpenStudy (anonymous):

It's C. x=3/2 or x=-3/2

OpenStudy (anonymous):

Yeah I was right

OpenStudy (anonymous):

4x² - 9 = 0 4x² - 9 = 0 write as difference of squares (2x - 3)² = 0 (2x)² - (3)² = 0 factor difference of squares 2x - 3 = 0 (2x - 3)(2x + 3) = 0 Set each factor = 0 & solve 2x = 3 2x - 3 = 0; 2x + 3 = 0 x = 3/2 2x = 3; 2x = -3 x = 3/2; x = -3/2 Correct solutions: 3/2 and -3/2

OpenStudy (anonymous):

thank youuu :))

OpenStudy (anonymous):

You'reee Welcome:)

OpenStudy (anonymous):

Your welcome